OpenStudy (anonymous):

(sqrt14 +sqrt10i)(14 - sqrt10i) and 3+ i/3-i

OpenStudy (eyust707):

\[(\sqrt{14} + \sqrt{10}i )(14 - \sqrt{10}i)\]

OpenStudy (eyust707):